The temperature T1 and T2 of the two heat reservoirs in an ideal carnotengine be 1500°C and 500°C respectively. which of these increasing T1 by100°C or decreasing T2 by 100°C would result in a greater improvementin the efficiency of the engine.

Grade:11

1 Answers

Arun
6 years ago
Dear Vaibhav
 
Efficiency = 1 - (T2/T1)
Case -1
n = (1873-773)/ 1873 = 58.73%
Case-2
n = (1773 - 673)/1773 = 62.04%
Clearly efficiency in 2nd case is higher.
 
Regards
Arun (askIITians forum expert)